Let X is Banach space and Let M be a subspace of X. Than M is itself a Banach space (using the norm from X) if and only if M is closed.
1
Expert's answer
2020-10-23T13:12:18-0400

1)Let "M" be a closed subset of "X", then "M=\overline{M}".

Take a fundamental sequence "\{a_n\}_{n\in\mathbb N}\subset M".

Since "M\subset X", we have that "\{a_n\}_{n\in\mathbb N}" is a fundamental sequence in "X". Then there is "\lim\limits_{n\to\infty}a_n=a" in "X".

Since "\{a_n\}_{n\in\mathbb N}\subset M", we have that "a\in\overline{M}=M" and so "\lim\limits_{n\to\infty}a_n=a" in "M".

By the definition of Banach space we have that "M" is a Banach space.

2)Let "M" be a Banach subspace of "X". Take arbitrary "a\in\overline{M}".

Then there is "\{a_n\}_{n\in\mathbb N}\subset M" such that "\lim\limits_{n\to\infty}a_n=a".

Since "\{a_n\}_{n\in\mathbb N}" is a convergent sequence, it is a fundamental sequence, so "\lim\limits_{n\to\infty}a_n=a\in M".

Since we take arbitrary "a\in\overline{M}", we have "\overline{M}\subset M".

We obtain "\overline{M}=M", because "M\subset\overline{M}", that is "M" is closed.